# Knowledge Transfer from Simple to Complex: A Safe and Efficient Reinforcement Learning Framework for Autonomous Driving Decision-Making

### v3

@@ 3.3 Simple to Complex Collaborative Decision-Making Framework
In this section, we introduce the S2CD framework in detail, which consists of an offline teacher training module, a high-level decision-making layer, and a low-level layer (comprising the path planning and control layers). The teacher training module first generates an autonomous driving teacher model by rapidly training a RL agent in a lightweight simulation environment. During the training of the teacher model, two neural networks are used to fit the Return function and the Q-Value function, respectively, predicting the return values and Q-values for state-action pairs.
The high-level decision-making layer utilizes guidance from the teacher model and state information from the perception module to generate high-level commands, which are then transmitted to the low-level layer. The low-level layer is responsible for managing the driving trajectory, including the driving path, lateral speed, and longitudinal speed. Ultimately, the ego vehicle executes actions based on the output values from the low-level layer. While our research primarily focuses on the high-level decision-making layer, the details of the low-level layer are provided in Section 3.3.3. Fig. 2 illustrates the components of the S2CD decision-making framework.
Figure 2: The S2CD decision-making framework consists of a teacher training module, a high-level decision-making layer, and a lower-level layer, enhanced by 4 innovative modules to improve learning efficiency and performance: 1. Employing the teacher model for action intervention and demonstration to enhance the safety of the student policy; 2. Utilizing dual-source data for training to improve sample efficiency; 3. Employing KL divergence constraints in policy updates to enable the student policy to quickly approach the teacher policy; 4. Gradually reducing the teacher model’s intervention via a weaning mechanism to prevent excessive reliance of the student on the teacher.
@@ 3.3.2 High-level Decision-Making Layer
First, after completing the offline training of the teacher model, the environment state $s_{t}$ is fed into the policy network $\pi^{t}$, which generates the probability distribution of the teacher’s policy and outputs the action $a^{t}_{t}$ with the highest probability ($a^{t}_{t}$ considered optimal by the teacher for $s_{t}$). Simultaneously, based on the Return network and Q-Value network, the return and Q-value for action $a^{t}_{t}$ in the given state $s_{t}$ are predicted. Next, the teacher’s guiding action $a^{t}_{t}$ is used as input to the student model, along with the state $s_{t}$, to form a new state representation $s_{t}^{\prime}$. The student agent’s policy network generates the probability distribution of actions based on $s_{t}^{\prime}$ and selects the action $a^{s}_{t}$ with the highest probability. At this stage, both the teacher and student models output actions $a^{t}_{t}$ and $a^{s}_{t}$, respectively, according to their own policies. To enhance the teacher’s guidance of the student and allow for intervention in the student’s actions when necessary, we propose a novel switch function based on the Q-values of $a^{t}_{t}$ and $a^{s}_{t}$, as shown below:
@@
The switch function indicates that the teacher will tolerate the student when the teacher cannot significantly outperform the student in terms of long-term returns. Specifically, when the difference in Q-value between actions $a^{t}_{t}$ and $a^{s}_{t}$ exceeds the tolerance threshold $(1-\tau)\varepsilon$, it implies that the student’s performance is suboptimal, necessitating intervention by the teacher. Otherwise, the teacher acknowledges the student’s performance and refrains from overriding the action. Here, Q-values are predicted by the trained Q-Value network, $\varepsilon$ is the tolerance coefficient used to control the extent of the student’s reliance on the teacher, and $\tau$ is a decay coefficient, where $1-\tau$ increases the teacher’s tolerance, gradually reducing the student’s dependence on the teacher. We design the decay coefficient $\tau$ as shown in Eq. 20:
@@
where $n_{e}$ is the number of episodes during training, $q_{1}$ and $q_{2}$ are hyperparameters shown in Table 1. Based on this switch function, the final action $a_{t}$, determined for execution, is passed to the low-level layer for vehicle control.
@@
Maximal learning rate 0.0005 Discount factor $\gamma$ 0.96
@@
Total steps of one episode 5,000 Clip parameter $\epsilon$ 0.2
@@
With the switch function, the return of the mixed behavior policy $J(\theta_{\text{{mix}}})$ is bounded both below and above by:
@@
Simultaneously, to illustrate the improvement in safety performance of S2CD resulting from teacher policy intervention, we introduce Theorem 4 to demonstrate that the expected cumulative reward obtained by policy $\pi^{\text{mix}}$ is guaranteed to be greater than or equal to the expected cumulative reward obtained by policy $\pi^{t}$, as expressed by:
@@
The expected cumulative reward obtained by learning from $\pi^{\text{mix}}$ is guaranteed to be greater than or equal to the expected cumulative reward obtained by $\pi^{t}$:
@@
When action $a_{t}$ is executed, the environment provides feedback, such as the return, to the student agent for updating the neural network. At this point, the student agent has two complete sets of data available for training and optimizing the policy $\pi^{s}$: $(s_{t}^{\prime},a_{t},r_{t})$ consists of on-policy data collected by executing action $a_{t}$ in the current state $s_{t}^{\prime}$ and receiving environmental feedback $r_{t})$, while the other set $(s_{t}^{\prime},a^{t}_{t},r^{t}_{t})$ is off-policy data provided by the teacher module, where action $a^{t}_{t}$ and the return $r^{t}_{t}$ are predicted based on state $s_{t}$. Due to the additional data provided by the teacher model, the S2CD framework achieves significantly higher sample efficiency. For the two distinct data sources from the teacher and student, the teacher’s guidance is particularly crucial, especially during the early stages of training. Therefore, to further enhance the learning efficiency of the algorithm, we applied the ACPPO algorithm proposed in Section 3.2 to optimize and update the neural network. ACPPO improves learning efficiency by loosening the update magnitude for significant samples while tightening it for less important ones.
@@
$L^{ACLIP^{\prime}}(\theta)=\left\{\begin{array}[]{ll}\mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\epsilon^{\prime}),1+(\epsilon-\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)-\xi C^{KL}_{t}\right],&{if}\ a_{t}=a^{s}_{t}\\[5.0pt] \mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\epsilon^{\prime}),1+(\epsilon+\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)-\xi C^{KL}_{t}\right],&{if}\ a_{t}=a^{t}_{t}\end{array}\right.$ (24)
@@
The teacher’s intervention can significantly accelerate the student’s learning process and prevent dangerous actions, thereby greatly enhancing the student’s model performance, especially during the early stages of training. However, since the teacher’s policy has an upper performance limit, over-reliance on the teacher’s guidance can constrain the student’s ultimate performance. To address this issue, we introduced a weaning mechanism that gradually reduces the teacher’s intervention in the student during the mid-to-late stages of training, thereby mitigating the influence of the teacher’s policy. This allows the S2CD framework to effectively balance teacher guidance with student exploration, ultimately achieving superior performance compared to a single policy. The weaning mechanism consists of two main components:
  • We introduce the decay factor $\tau$, which gradually increases the tolerance of the teacher’s policy toward the student. Specifically, the tolerance increases by $1-\tau$, leading to a progressive reduction in the intervention probability $\omega$ of the teacher’s actions, ultimately decreasing it to 0.
@@
$L^{S2CD}(\theta)=\left\{\begin{array}[]{ll}\mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\tau\epsilon^{\prime}),1+(\epsilon-\tau\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)-\tau\xi C^{KL}_{t}\right],&{if}\ a_{t}=a^{s}_{t}\\[5.0pt] \mathbb{E}_{\pi_{\theta}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\tau\epsilon^{\prime}),1+(\epsilon+\tau\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a)\right)-\tau\xi C^{KL}_{t}\right],&{if}\ a_{t}=a^{t}_{t}\end{array}\right.$ (25)
According to the decay coefficient, in the early stages of training, as much prior knowledge as possible is provided to guide ACPPO learning. In the middle and late stages of training, it becomes necessary to gradually reduce the influence of the teacher policy on the student. And as decay coefficient gradually decreases, the objective function $L^{S2CD}(\theta)$ will regress to the objective function of the regular PPO algorithm, as shown in Eq. 11. We update the ACPPO agent by maximizing $L^{S2CD}(\theta)$.
The process of the S2CD decision-making framework is outlined in Algorithm 2, and all parameters are listed in Table 1.
@@ 3.3.3 Low-level Layer
In our proposed S2CD decision-making framework, we have designed a hierarchical architecture to facilitate the smooth driving of autonomous vehicles in complex and dynamic traffic environments. Initially, the high-level decision-making layer generates either a follow or a lane-change command based on the state $s_{t}^{\prime}$. Subsequently, the low-level layer executes path planning and controls the vehicle’s driving trajectory and speed in accordance with the high-level commands. As this paper primarily concentrates on high-level decision-making algorithms, we have opted for classic and reliable algorithms for the low-level layer.
@@
In the path planning layer, we employ cubic spline functions to plan the vehicle’s driving path. The centerline points of the driving path are defined as $(x_{0},y_{0}),(x_{1},y_{1}),\dots,(x_{n},y_{n})$. Each $S_{i}$, representing the segment along the interpolated centerline, is bounded by $(x_{i},y_{i})$ and $(x_{i+1},y_{i+1})$. The cubic spline function is defined as follows:
@@
where parameters $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$ can be solved using the methods outlined in Paper Jiang et al. (2020).
The global trajectory is determined by Carla’s navigation points. When the high-level command is a lane change, we set the lateral position of the next navigation point on the centerline of the target lane, maintaining a longitudinal distance of 10 m from the current position. Subsequently, cubic spline interpolation is employed for local path planning along this segment of the trajectory.
@@
In this paper, we exclusively focus on lateral decision-making, which encompasses lane changes as well as following behavior. To maintain simplicity and prevent the over-conservative behavior of the agent, acceleration and speed are controlled by the Intelligent Driver Model (IDM) (Treiber et al., 2000). The target acceleration $\dot{v}_{e}$ of the ego vehicle can be calculated using Eq. 27:
@@
where $v_{e}$ is the speed of the ego vehicle , $v_{1}$ is the speed of the front vehicle, and other parameters of IDM are detailed in Table 3. With the target acceleration and current speed, the next moment’s target speed can be determined.
@@
Once the reference driving path and target speed are obtained, the control layer utilizes the PID algorithm to calculate the values for the vehicle’s throttle, brake, and steering angle, employing two sets of PID parameters separately for lateral control and longitudinal control. All PID parameters are shown in Table 3.
@@ 4 Implementation
This section details the implementation of our simulation experiments. First, we compare the differences between the Highway-Env and Carla simulation environments. Then, we provide a detailed description of the medium-density highway lane-change scenario. Next, we model the driving scenario as a Markov Decision Process (MDP). Finally, we specify the parameter settings used in the simulation.
@@ 4.2 Scenarios Specification
Medium traffic density on highways is characterized by approximately 15 vehicles per lane per kilometer, with vehicle spacing ranging from 50 m to 90 m (Gold et al., 2016). This density strikes a balance by avoiding congested traffic, which hinders lane changes, and low-density traffic, where lane changes become unnecessary. Consequently, it serves as a typical representation of most lane-change scenarios. We employ the medium-density traffic scenario as a comprehensive platform to train and evaluate the performance of the S2CD lane-change decision-making framework. Fig. 3 illustrates the scenario’s performance in the Highway-Env and Carla simulation environments, where the ego vehicle navigates a 3-lane highway and executes lane changes while surrounded by other vehicles. Consistent with driving knowledge and applicable traffic regulations, the ego vehicle must consider the positions and movements of vehicles ahead, as well as those to its front-left, rear-left, front-right, and rear-right, during lane-change decision-making. Therefore, the real-time states of these surrounding vehicles (such as speed and distance) are integrated into the model’s input.
In the experiment, both the ego vehicle and surrounding vehicles start with an initial speed of 0 m/s, while the ego vehicle’s maximum speed is limited to 25 m/s. To more accurately replicate real-world driving conditions, the target speeds of the surrounding vehicles are randomly selected within the range of 15 to 25 m/s. Poor lane-change decisions may result in collisions between the ego vehicle, other vehicles, or road boundaries. An episode is considered complete if the ego vehicle successfully drives 1 km without any collisions. If a collision occurs, the episode is marked as failed, and the environment is reset for the next episode.

### v4

@@ 3.3 Simple to Complex Collaborative Decision-Making Framework
In this section, we provide a detailed overview of the components of the S2CD framework, which include the offline teacher training module, the high-level layer, and the low-level layer. In the teacher training module, an RL agent is rapidly trained in a lightweight simulation environment as the teacher model, with two neural networks used to approximate the Return function and the Q-Value function, respectively, to predict the return values and Q-values for state-action pairs. The decision-making layer outputs final high-level commands based on the teacher’s action guidance and state information, while the low-level layer receives these commands to perform path planning and ultimately control the vehicle’s motion. Fig. 1 depicts the components of the S2CD framework.
Figure 1: The S2CD framework consists of a teacher training module, a high-level decision-making layer, and a lower-level layer, enhanced by 4 innovative modules to improve learning efficiency and performance: 1. Employing the teacher model for action intervention and demonstration to enhance the safety of the student policy; 2. Utilizing dual-source data for training to improve sample efficiency; 3. Employing KL divergence constraints in policy updates to enable the student policy to quickly approach the teacher policy; 4. Gradually reducing the teacher model’s intervention via a weaning mechanism to prevent excessive reliance of the student on the teacher.
@@ 3.3.2 High-level Decision-Making Layer
First, after completing the offline training of the teacher model, the environment state $s_{t}$ is fed into the teacher’s policy network $\pi^{t}$, which outputs the action $a^{t}_{t}$ with the highest probability ($a^{t}_{t}$ considered optimal by the teacher for $s_{t}$). Simultaneously, based on the Return network and Q-Value network, the return and Q-value for action $a^{t}_{t}$ in the given state $s_{t}$ are predicted. Next, the teacher’s guiding action $a^{t}_{t}$ is used as input to the student model, along with the state $s_{t}$, to form a new state representation $s_{t}^{\prime}$. The student agent’s policy network generates the probability distribution of actions based on $s_{t}^{\prime}$ and selects the action $a^{s}_{t}$ with the highest probability. At this stage, both the teacher and student models output actions $a^{t}_{t}$ and $a^{s}_{t}$, respectively, according to their own policies. To enhance the teacher’s guidance of the student and allow for intervention in the student’s actions when necessary, we propose a novel switch function based on the Q-values of $a^{t}_{t}$ and $a^{s}_{t}$, as shown below:
@@
The switch function indicates that the teacher will tolerate the student when the teacher cannot significantly outperform the student in terms of long-term returns. Specifically, when the difference in Q-value between actions $a^{t}_{t}$ and $a^{s}_{t}$ exceeds the tolerance threshold $(1-\tau)\varepsilon$, it implies that the student’s performance is suboptimal, necessitating intervention by the teacher. Otherwise, the teacher acknowledges the student’s performance and refrains from overriding the action. Here, Q-values are predicted by the trained Q-Value network, $\varepsilon$ is the tolerance coefficient used to control the extent of the student’s reliance on the teacher, and $\tau$ is a decay coefficient, where $1-\tau$ increases the teacher’s tolerance, gradually reducing the student’s dependence on the teacher. $\tau$ is shown in Eq. 17:
@@
where $n_{e}$ is the number of episodes experienced, and the values of $q_{1}$ and $q_{2}$ are shown in Table 1. Based on this switch function, the final action $a_{t}$ can be determined for execution.
@@
Discount factor $\gamma$ 0.96 Learning rate 0.0005
@@
Clip parameter $\epsilon$ 0.2 Total steps per episode 5,000
@@
With the switch function, the return $J(\theta_{\text{{mix}}})$ is bounded both below and above by:
@@
Simultaneously, to illustrate the improvement in safety performance of S2CD resulting from teacher policy intervention, we introduce Theorem 2 to demonstrate that the expected cumulative reward obtained by policy $\pi^{\text{mix}}$ is assured to be no less than the expected cumulative return gained by policy $\pi^{t}$, as expressed by:
@@
The expected cumulative reward acquired from $\pi^{\text{mix}}$ is ensured to be no less than the expected cumulative reward derived from $\pi^{t}$:
@@
When action $a_{t}$ is executed, the environment provides feedback, such as the return, to the student agent for updating the neural network. At this point, the student agent has two complete sets of data available for training and optimizing the policy $\pi^{s}$: $(s_{t}^{\prime},a_{t},r_{t})$ consists of on-policy data collected by executing action $a_{t}$ in the current state $s_{t}^{\prime}$ and receiving environmental feedback $r_{t})$, while the other set $(s_{t}^{\prime},a^{t}_{t},r^{t}_{t})$ is off-policy data provided by the teacher module, where action $a^{t}_{t}$ and the return $r^{t}_{t}$ are predicted based on state $s_{t}$. Due to the additional data provided by the teacher model, the S2CD framework achieves significantly higher sample efficiency. For the two distinct data sources from the teacher and student, the teacher’s guidance is particularly crucial, particularly in the initial phases of training. Therefore, we applied the ACPPO+ algorithm proposed in Section 3.2 to optimize and update the neural network, to further enhance the learning efficiency of the algorithm. ACPPO+ improves learning efficiency by loosening the update magnitude for significant samples while tightening it for less important ones.
@@
$\begin{split}L^{ACLIP+^{\prime}}(\theta)&=\mathbb{E}_{\pi_{\theta}}\Bigg{[}\mathbb{I}_{a_{t}^{s}}\cdot\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a_{t}^{s}),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\epsilon^{\prime}),1+(\epsilon-\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a_{t}^{s})\right)\\ &\quad+\mathbb{I}_{a_{t}^{t}}\cdot\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a_{t}^{t}),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\epsilon^{\prime}),1+(\epsilon+\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a_{t}^{t})\right)-\xi C^{KL}_{t}\Bigg{]}\end{split}$ (21)
@@
The teacher’s intervention can significantly accelerate the student’s learning process and prevent dangerous actions, thereby greatly enhancing the student’s model performance, especially during the early stages of training. However, since the teacher’s policy has an upper performance limit, over-reliance on the teacher’s guidance can constrain the student’s ultimate performance. To address this issue, we introduced a weaning mechanism that gradually reduces the teacher’s intervention in the student during the mid-to-late stages of training, thereby mitigating the influence of the teacher’s policy. This allows the S2CD framework to effectively balance teacher guidance with student exploration, ultimately achieving superior performance compared to a single policy. The weaning mechanism is composed of two main components:
  • We use the decay coefficient $\tau$, which gradually increases the tolerance of the teacher’s policy toward the student. Specifically, the tolerance increases by $1-\tau$, leading to a progressive reduction in the intervention probability $\omega$ of the teacher’s actions, ultimately decreasing it to 0.
@@
$\begin{split}L^{S2CD}(\theta)&=\mathbb{E}_{\pi_{\theta}}\Bigg{[}\mathbb{I}_{a_{t}^{s}}\cdot\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a_{t}^{s}),\text{clip}\left(r_{t}(\theta),1-(\epsilon+\tau\epsilon^{\prime}),1+(\epsilon-\tau\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a_{t}^{s})\right)\\ &\quad+\mathbb{I}_{a_{t}^{t}}\cdot\min\left(r_{t}(\theta)A^{\pi_{\theta}}(s,a_{t}^{t}),\text{clip}\left(r_{t}(\theta),1-(\epsilon-\tau\epsilon^{\prime}),1+(\epsilon+\tau\epsilon^{\prime})\right)A^{\pi_{\theta}}(s,a_{t}^{t})\right)-\tau\xi C^{KL}_{t}\Bigg{]}\end{split}$ (22)
According to $\tau$, we allow the teacher to provide as many correct decisions as possible during the early stages of training to intervene and guide the student’s learning. As training progresses, teacher intervention is gradually reduced in the later stages, allowing the student to independently explore better strategies. Eventually, as $\tau$ decreases and approaches zero, $L^{S2CD}(\theta)$ gradually returns to the form of Eq. 11. The operation of the S2CD decision-making framework is shown in Algorithm 2, and all parameters are listed in Table 1.
@@ 3.3.3 Low-level Layer
In the S2CD framework, we developed a hierarchical architecture to enable autonomous vehicles to drive smoothly. This framework consists of two components: high-level decision-making and low-level execution. First, the high-level layer outputs lane-change or car-following commands according to the current state, providing a macro driving strategy for the vehicle. Next, the low-level execution layer receives these high-level commands and is responsible for path planning, trajectory and speed control to precisely carry out the selected driving strategy.
@@
We represent the centerline of the path with a series of points $(x_{0},y_{0}),(x_{1},y_{1}),\dots,(x_{n},y_{n})$, where each segment $S_{i}$ represents a portion along the centerline, bounded by points $(x_{i},y_{i})$ and $(x_{i+1},y_{i+1})$. The vehicle’s driving path is then planned using a cubic spline function, defined as follows:
@@
where $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$ are parameters determined by the approach outlined in Jiang et al. (2020).
The ego vehicle’s global trajectory is determined by the built-in path navigation of the Carla simulator. When the decision-making layer issues a lane-change command, the global navigation point is positioned in the target lane at a distance of 10 meters from the vehicle’s current position. Eq. 23 is then used to determine the trajectory of the lane-change segment for local path planning.
@@
Since this study focuses on lane-change and car-following decision-making algorithms, the acceleration and speed of the ego vehicle are controlled using the Intelligent Driver Model (IDM) (Treiber et al., 2000) to ensure simplicity and prevent overly conservative behavior by the agent. Therefore, the target acceleration $\dot{v}_{e}$ is calculated using Eq. 24.
@@
where $v_{e}$ and $v_{1}$ represent the speeds of the ego and the lead vehicles, respectively, while other parameters in the 24 are detailed in Table 3. Using $v_{e}$ and $\dot{v}_{e}$, the target speed for the next moment can be calculated.
After establishing the target speed and driving path, the PID algorithm is applied to determine the specific control values. To enhance control performance, two sets of parameters are utilized, as detailed in Table 3.
@@
Target speed $v_{0}$ 25 m/s
Target time gap $T$ 0.6 s
Safety distance $s_{0}$ 2.0 m
@@ 4 Implementation
This section begins by comparing the differences between the Highway-Env and Carla simulation environments, followed by a detailed description of a medium-density highway lane-change scenario. Next, the scenario is modeled as a Markov Decision Process (MDP), and finally, all settings used in the simulation experiments are introduced.
@@ 4.2 Scenarios Specification
Medium traffic density on highways is defined as having roughly 15 vehicles per lane per kilometer, with vehicle spacing between 50 and 90 meters (Gold et al., 2016). This density level provides a balance by avoiding both congested conditions, which make lane changes difficult, and low-density conditions, where lane changes are largely unnecessary. Therefore, it represents a typical scenario for most lane-change situations. We use this scenario as the platform for training and evaluating the performance of the S2CD framework. Fig. 2 shows the traffic scenario on the three-lane highway in the Highway-Env and Carla simulation environments, where the ego vehicle performs a lane change while surrounded by other vehicles. According to driving principles and relevant traffic regulations, the ego vehicle must consider the positions and movements of the front, front-left, rear-left, front-right, and rear-right vehicles when executing a lane change. Therefore, the states of the ego vehicle and surrounding vehicles, including their speeds and distances, will serve as inputs to the model.
In the experiments, all vehicles start with an initial velocity of 0 m/s, with a maximum velocity of 25 m/s. To more accurately simulate real traffic conditions, the target speeds of vehicles are set between 15 and 25 m/s. If the ego vehicle makes improper lane-change decisions while driving, it may come too close to other vehicles or even collide with them or the road boundary. If the ego vehicle successfully travels 1 kilometer without any collisions, the episode is considered successful. In the event of a collision, the episode is marked as a failure, and the environment is immediately reset for the next episode.